# TESGNN: Temporal Equivariant Scene Graph Neural Networks for Efficient and Robust Multi-View 3D Scene Understanding

### v1

@@ 6.1 Dataset and Metrics
Additional metrics: We assess our Temporal Model by following the Information Retrieval evaluation metrics, including R@k, mAP@k, and MRR@k. mAP@k measures the quality of rankings by emphasizing $k$ relevant nodes at higher ranks. MRR@k evaluates the effectiveness of retrieval by averaging the reciprocal ranks of the $k$ similar nodes. We also report the resource used for training and inference.
6.2 Scene Graph Generation Evaluation
Superior prediction performance: Table 1 compares the results between ESGNN (ours) with existing models 3DSSG [52] and SGFN [53] on the 3DSSG-l20 dataset with geometric segmentation setting. Our method obtains high results in both relationship, object, and predicate classification. Especially, ESGNN outperforms the existing methods in relationship prediction and obtains significantly higher R@1 in predicate compared to SGFN. ESGNN also works well with unseen data, with competitive result compared to SGFN, as shown in Table 2.
@@ 6.2 Scene Graph Generation Evaluation
Table 1: Scene graph predictions for relationship triplet, object, and predicate, measured on 3DSSG-l20. The Recall column reports the recall scores on objects (Obj.) and relationships (Rel.)
@@
Table 2: Scene graph predictions for new unseen relationship triplet, object, and predicate, measured on 3DSSG-l20 with geometric segmentation.
Faster training convergence: Figure 5(a) and 5(d) report the recalls for nodes (objects) and edges (relationships) during training between ESGNN and SGFN on both train and validation sets. The recall slope of ESGNN in the first 10 epochs (5000 steps) is significantly higher than that of SGFN. This indicates that ESGNN has faster convergence and higher initial recall.
@@
Figure 6: Edge and node confusion matrices after the first 1000 steps (2 epochs). ESGNN is better at handling imbalanced classes, as the predictions spread across multiple classes rather than “blindly” predicting towards dominant classes like SGFN.
@@
6.3 Ablation Study
Table 3 evaluates of ESGNN with different architectures and settings that we experimented. ⓪ is the SGFN, run as the baseline model for comparison. ① is the ESGNN with 1 FAN-GCL layer and 1 EGCL layer. This is our best performer and is used for experiments in Section 6.2. ② is similar to ①. The only difference is that we concatenate coordinate embedding to the output edge embedding after message passing. Meanwhile, ③ and ④ are similar to ① and ② respectively, but with 2 FAN-GCL layers and 2 EGCL layers.
@@ 6.3 Ablation Study
Table 3: Evaluation of different ESGNN architectures on scene graph generation task on 3DSSG-l20 dataset. ② is our best performer and is used for the evaluation in Section 6.2.
@@
Models ③ and ④ perform well on the train set but poorly on the validation and test sets, potentially suffering overfitting as they contain more layers. Models ② and ④ have higher edge recalls in several initial epochs, but experience a decline in the later epochs, as shown in Fig 5(b), 5(e).
6.4 ESGNN with Image Encoder
Our model also poses a potential in application on point-image encoders model together such as JointSSG [54]. We implement our GNN architecture similar to JointSSG and name it Joint-ESGNN. Figures 5(c) and 5(f) show the performance of our model with image encoders compared to JointSSG and SGFN.
6.5 Temporal Graph Matching Evaluation
[…]
Metric@k k=1 k=2 k=3 k=5
R@k 0.7015 0.8406 0.9009 0.9525
mAP@k 0.7015 0.7893 0.8194 0.8409
MRR@k 0.7015 0.7711 0.7911 0.8032
@@ 6.5 Temporal Graph Matching Evaluation
Table 4: Evaluation for our Temporal Model with 10 training epochs. Notice that k=1 is the accuracy of the node-matching.
Prediction performance: Our Temporal Model is trained on 321,344 pairs of node with 16,161 positive pairs generated from the 3DSSG dataset. Table 4 demonstrates its strong accuracy in node-matching, highlighting its ability to effectively identify similar nodes.
Potential Customization and Improvement: As graph-matching for point cloud scene graphs is a novel approach, no existing methods are available for direct comparison. Given our model’s embedding-based nature, there are various opportunities for improvement. For instance, incorporating multiple loss functions, such as the Multiple Negatives Ranking Loss [13] alongside Contrastive Loss, could significantly enhance training performance. Various data sampling strategies could also be explored to enhance the model’s robustness.
6.6 Run time
We measure the runtime of our system during the training phase, which demands the most resources. Our machine is equipped with two Intel Xeon Gold 6242 CPUs, each running at 3.9 GHz with a total of 64 threads, and four Nvidia GeForce RTX 3090 GPUs. However, we utilized only 8 data workers during training, resulting in a total CPU utilization of 0.3%. Only one GPU was used, consuming a maximum of 1.6 GB of memory per experiment. The average training time per step, excluding data loading, is approximately 0.25 seconds. Inference time is significantly shorter, and our model’s size is only 37 MB.
7 Conclusion
@@
In this paper, we introduced the TESGNN, a novel method that overcomes key limitations in 3D scene understanding. By leveraging the symmetry-preserving property of the Equivariant GNN, our architecture ensures robust and efficient generation of semantic scene graphs from 3D point clouds. Our proposed Temporal Model integrates multiple scene graphs, providing a global representations for real-time dynamic environments. Experimental results show that TESGNN outperforms state-of-the-art methods in accuracy, convergence speed, and computational efficiency.
Limitations & Future Work
@@
Future work will focus on optimizing TESGNN for complex real-world scenarios by integrating additional sensor modalities like LiDAR and RGB-D data to improve scene graph generation. We also aim to extend the model’s temporal capabilities for continuous data streams, enhancing its suitability for autonomous navigation and multi-agent systems.

### v2

@@ VI-B Metrics
[…]
VI-C Scene Graph Extraction Evaluation
Better prediction outcome: Table I compares the results between ESGNN (ours) with existing models 3DSSG [10] and SGFN [11] on the 3DSSG-l20 dataset with geometric segmentation setting. Our method obtains high results in both relationship, object, and predicate classification. Especially, ESGNN outperforms the existing methods in relationship prediction and obtains significantly higher R@1 in predicate compared to SGFN. For unseen data, ESGNN is competitive with SGFN, as shown in Table II.
@@ VI-C Scene Graph Extraction Evaluation
TABLE I: Scene graph predictions for relationship triplet, object, and predicate, measured on 3DSSG-l20. The Recall column reports the recall scores on objects (Obj.) and relationships (Rel.)
@@
TABLE II: Scene graph predictions for new unseen relationship triplet, object, and predicate, measured on 3DSSG-l20 with geometric segmentation.
Faster training convergence: Figure 5(a) and 5(d) report the recalls for nodes (objects) and edges (relationships) during training of ESGNN and SGFN on both train and validation sets. The recall slope of ESGNN in the first 10 epochs (5000 steps) is significantly higher than that of SGFN. This shows that ESGNN has faster convergence and higher initial recall.
@@
Figure 6: Confusion matrices after the first 1000 steps (2 epochs). ESGNN is better at handling imbalanced classes, as the predictions spread across multiple classes rather than “blindly” predicting towards dominant classes like SGFN.
@@
Scene Graph Extractor with Image Encoder: ESGNN are also potential to apply with point-image encoder instead of 3D point clouds only. Figures 5(c) and 5(f) show the performance of our implementation with image encoders, called Joint-ESGNN, compared to existing methods including JointSSG [12] and SGFN.
VI-D Ablation Study
Table III evaluates of ESGNN with different architectures and settings that we experimented. ⓪ is the SGFN, run as the baseline model for comparison. ① is the ESGNN with 1 FAN-GCL layer and 1 EGCL layer. This is our best performer and is used for experiments in Section VI-C. ② is similar to ①. The only difference is that we concatenate coordinate embedding to the output edge embedding after message passing. ③ and ④ are similar to ① and ② respectively, but with 2 FAN-GCL layers and 2 EGCL layers.
@@ VI-D Ablation Study
TABLE III: Evaluation of different ESGNN architectures on scene graph generation task on 3DSSG-l20 dataset. ② is our best performer and is used for evaluation in Section VI-C.
@@
Models ③ and ④ perform well on the train set but poorly on the validation and test sets, potentially suffering overfitting as they contain more layers. Models ② and ④ have higher edge recalls in several initial epochs, but decline in the later epochs, as shown in Fig 5(b), 5(e).
VI-E Temporal Graph Matching Evaluation
Experiment Setup: We assess our work from two key perspectives. First, we evaluate our Graph Matching Network using different backbone scene graph extractors, including the state-of-the-art SGFN [11] and our proposed ESGNN. Second, we compare the effectiveness of the SGFN and ESGNN backbones on the state-of-the-art SG-PGM [30]. To ensure the same set-up, we reimplement and evaluate SG-PGM with our proposed dataset mentioned in Section VI-A.
Method Backbone R@k MRR@k
@1 @2 @3 @5 @2 @3
SG-PGM SGFN 10.61 19.16 28.56 44.71 14.88 18.02
ESGNN $\mathbf{30.14}$ $\mathbf{46.94}$ $\mathbf{57.63}$ $\mathbf{73.86}$ $\mathbf{38.54}$ $\mathbf{42.10}$
SG-PGM SGFN 18.54 33.28 $\mathbf{45.73}$ $\mathbf{62.82}$ 25.91 30.06
(point) ESGNN $\mathbf{20.91}$ $\mathbf{33.55}$ 44.41 62.01 $\mathbf{27.23}$ $\mathbf{30.85}$
Ours SGFN 54.48 72.95 82.11 90.45 63.71 66.77
ESGNN $\mathbf{70.15}$ $\mathbf{84.06}$ $\mathbf{90.09}$ $\mathbf{95.25}$ $\mathbf{78.93}$ $\mathbf{81.94}$
@@ VI-E Temporal Graph Matching Evaluation
TABLE IV: Evaluation for Graph Matching Network with 30 training epochs. Notice that k=1 is the accuracy of the node-matching and R@1 = MRR@1
Better outcome with ESGNN backbone scene graph extractor: Table IV demonstrates that utilizing our proposed ESGNN as the backbone scene graph extractor consistently outperforms SGFN across all metrics, on both graph matching networks. This underscores the advantage of ESGNN’s symmetry-preserving property, which enhances the representation of nodes and edges in 3D point clouds, making them more resilient to transformations and significantly improving graph matching performance.
Better outcome with our Graph Matching Network: Although SG-PGM [30] achieved competitive results for node matching in their setting, it performs poorly in our scenario, where node and edge features are extracted from a scene graph extractor model rather than one-hot encoded from the ground truth. In our case, both nodes and edges have more labels, leading to several ambiguous predictions, including overlaps, incorrect assignments, and permutations. This explains why SG-PGM fails in this setting. In contrast, our model leverages the equivariant properties of nodes and edges while applying sum pooling, ensuring permutation invariance and significantly outperforming SG-PGM in this scenario, indicating the robustness of the model.
VII Conclusion
@@
In this paper, we introduced the TESGNN, a novel method that overcomes key limitations in 3D scene understanding. By leveraging the symmetry-preserving property of the Equivariant GNN, our architecture ensures robust and efficient generation of semantic scene graphs from 3D point clouds. Our proposed Temporal Graph Matching Model provides a global representation from local scene graphs for real-time dynamic environments. Experimental results show that TESGNN outperforms state-of-the-art methods in accuracy, convergence speed, and computational efficiency.
@@
Limitations & Future Work Future work will focus on optimizing TESGNN for complex real-world scenarios by integrating additional sensor modalities like LiDAR and RGB-D data to improve scene graph generation. We also aim to extend the model’s temporal capabilities for continuous data streams, enhancing its suitability for autonomous navigation and multi-agent systems.